### v1

@@ 5.1 Service Placement Policy
The provisioner component serves as the foundation for other policies in CloudNativeSim, providing the mapping of various resources such as CPU cycles, memory, and storage. During the simulation runtime, the provisioner component continuously calculates the total and available resources in real-time, as well as which entities are utilizing these resources.
@@
6:        if $instance.require\_resources\leq vm.idel\_resources$ then
@@
5.2 Usage History Upadating
To observe the state of a microservice-based system and to guide the simulator’s subsequent scheduling activities, a model of resource utilization is essential. In CloudSim, the utilization model serves a similar function, but its inputs only consist of time, and it is dynamically changing based on extensive historical data from real machines, such as those from PlanetLab 25. While this approach is applicable to cloud-native applications, it lacks sufficient interpretability and struggles with the challenges posed by dynamic environments. Consequently, we propose a more dynamic method that updates the Usage History to record the resource status within the cloud-native application during simulation runtime.
@@ 5.2 Usage History Upadating
Based on above observations, we attempt to quantify resource usage by the number and duration of cloudlets assigned to an instance for simulation. To address this, CloudNativeSim employs a linear approach to correlate the performance of instances with cloudlets’ resource usage. Using the specific calculation method, a cloudlet’s execution time depends on the CPU time allocated to it and the processing speed of the instance executing it. The execution time of cloudlets ultimately affects the calculation of response time. Similarly, the memory usage and bandwidth consumption of cloudlets are calculated using a similar approach. Once the resource consumption by the cloudlets is calculated, the corresponding instance’s usage data is updated and recorded in the usage history.
@@ 5.3 Service Scaling Policy
In the simulation of microservice-based architectures, the service scaling policy is a core function. It can leverage historical resource utilization data to maintain optimal performance, while preventing both resource wastage and inefficiency 26. During the operation of the simulator, a service scaling event is triggered at regular intervals to automatically check whether the current service requires scaling, such as when an instance’s CPU utilization consistently remains high 27. Once scaling is deemed necessary, the user-defined scaling plan is executed. In CloudNativeSim, we provide the most common horizontal and vertical scaling solutions, as outlined in Algorithms 5.3 and 5.3.
@@
Utilizing scaling policies greatly enhances the dynamism of CloudNativeSim, enabling it to flexibly respond to changes in workload, ensuring high service availability and efficient resource utilization. In Section 6.4, we will discuss the impacts of different scaling policies. Users can also customize auto-scaling policies, such as those based on response time or SLO violation rates 28, workload levels 29, and may even implement hybrid policies 30, 31.
In this section, we provide a detailed overview of CloudNativeSim’s service scaling policies. By periodically triggering service scaling events, the simulator can automatically detect whether the current service requires scaling or not. Implementing these scaling policies significantly enhances the dynamism of CloudNativeSim, enabling it to flexibly respond to workload changes, ensuring high service availability and efficient resource utilization.
@@ 6.1 Tests of Simulation Capacity
To test the simulation capacity of CloudNativeSim, we conducted a capacity test on a common PC with a 16-core, 2.2 GHz CPU, and limited the running memory to 4 GB. To evaluate the simulator’s extreme capacity, we considered the following parameters: the number of services, the total number of instances, the number of requests, and the number of cloudlets. Various parameter sets resulted in different running times on the realistic testbed (in seconds).
To ensure the reliability of our results, we designed 4 cases and tested 2 parameter sets for each case, conducting 5 repetitive tests for each set. Table 2 presents the average running times for these tests:
@@
To avoid randomness, we conducted ablation experiments in Case 1 and Case 2, selecting parameter sets with running times under 15 seconds for analysis. In Case 1, we limited the number of services to 1 and the number of instances to one thousand. Under these conditions, CloudNativeSim can simulate up to $10^{6}$ requests. In Case 2, we fixed the number of instances at 1 and the number of requests at $10^{4}$, demonstrating that CloudNativeSim can handle over $5\times 10^{4}$ services. Notably, the number of cloudlets also increases with the number of requests and services, leading to a substantial rise in the simulator’s internal running time.
@@ 6.2 Verification of Request Generation
To validate the performance of the Request Generator and the correctness of modeling in Section 4.3.1, we select the most important parameters $(N_{c},v,p)$ to conduct four sets of test experiments on the Request Generator. By collecting and plotting the growth curves of number of clients, QPS, and total requests, we aimed to evaluate the performance of the Request Generator. Figure 9 illustrates the results produced by the Request Generator across four different test configurations:
@@
In Figure 9, the segmented curves in the three subplots exhibit distinct characteristics around 100 seconds, coinciding with the period of customer number growth, $\frac{N_{c}}{v}$. This period is highlighted due to its significance. During this time, the number of clients grows linearly at the spawn rate $v$, eventually stabilizing at the client number $N_{c}$. This process remains independent of the waiting interval $p$, which explains the presence of only three curves. This behavior aligns with the description provided in Equation 1. Initially, the QPS shows approximately linear growth before reaching an oscillatory convergence state. The speed of convergence is influenced by the spawn rate $v$, while the final converged value is determined by both $N_{c}$ and $p$. Specifically, the final value is negatively correlated with the value and interval size of $p$ and positively correlated with $N_{c}$, consistent with the description in Equation 3. In subplot 9(b), the blue and orange curves, which share the same $p$ and $N_{c}$, converge closely and oscillate around 100, as derived from Equation 3. For the total number of requests, the curve exhibits a nonlinear growth trend in the first segment and a linear growth trend in the second segment, consistent with the description in Equation 4.
@@ 6.3 Evaluation of Response Time Simulation
To assess the simulator’s capability in approximating the response time of actual applications, we conducted a stress test on the cluster using Locust, collecting average response times under varying numbers of users as the reference standard. In the Locust script, user wait times were set between 5 to 15 seconds, with a duration limit of 600 seconds. The same settings were applied in the simulator.
@@ 6.4 Demonstration of Scaling Policies
7 Conclusions and Future Work
@@
There are a number of future directions that can enhance CloudNativeSim capabilities and resource management strategies:
  • Modeling of hybrid services and serverless environments. Hybrid services and serverless environments are emerging fields and important components of cloud-native environments. Modeling these scenarios can further enhance the utility of CloudNativeSim, enabling it to support a broader range of use cases.
  • Support for various tasks. Considering that tasks in cloud-native environments are various (e.g., data-intensive tasks), each type requires different resource consumption models. By supporting a wider variety of tasks, CloudNativeSim can improve its simulation performance and provide more accurate resource management insights.
  • Developing the graphical user interface (GUI). A more interactive and user-friendly GUI can significantly improve the user experience. An intuitive interface will make CloudNativeSim more accessible to a broader audience, including users with less technical expertise, thereby increasing its adoption and usability.
  • Refining network traffic modeling. Accurate network traffic modeling is crucial for understanding the performance and scalability of microservices, as network overhead can significantly impact overall system performance. Enhancing the simulation of bandwidth utilization during data transmission and reception between microservice entities will provide more realistic and valuable simulations, aiding in the design and optimization of cloud-native applications.
  • Optimizing large-scale scenario simulations. As cloud-native applications continue to grow in scale and complexity, the ability to simulate extensive and intricate scenarios becomes essential. To meet this challenge, CloudNativeSim needs to implement better memory management practices, such as removing unused objects and using more efficient data structures and loop statements. These improvements will enhance its capability to handle large-scale simulations efficiently.
@@ Acknowledgments
This work is supported by National Key R & D Program of China (No.2021YFB3300200), National Natural Science Foundation of China (No. 62072451, 62102408, 92267105), Guangdong Basic and Applied Basic Research Foundation (No. 2023B1515130002, 2024A1515010251), Guangdong Special Support Plan (No. 2021TQ06X990), Shenzhen Science and Technology Program (Grant No. RCBS20210609104609044), Shenzhen Basic Research Program (No. JCYJ20220818101610023, KJZD20230923113800001), and Shenzhen Industrial Application Projects of undertaking the National key R & D Program of China (No.CJGJZD20210408091600002).
The authors also thank Dr. Kangjin Wang from Peking University (developer of ColocationSim) and Dr. Xiang He from Harbin Institute of Technology (developer of ServiceSim) for their valuable suggestions on improving the design and implementation of CloudNativeSim.

### v2

@@ 5.1 Service Placement Policy
The provisioner component, which manages and tracks resource allocation and utilization, serves as the foundation for other policies in CloudNativeSim, providing the mapping of various resources such as CPU cycles, memory, and storage. During the simulation runtime, the provisioner component continuously calculates the total and available resources in real-time, as well as which entities are utilizing these resources.
@@
6:        if $instance.require\_resources\leq vm.idle\_resources$ then
@@
5.2 Usage History Updating
To observe the state of a microservice-based system and to guide the simulator’s subsequent scheduling activities, a model of resource utilization is essential. In CloudSim, the utilization model serves a similar function, but its inputs only consist of time, and it is dynamically changing based on extensive historical data from real machines, such as those from PlanetLab 26. While this approach is applicable to cloud-native applications, it lacks sufficient interpretability and struggles with the challenges posed by dynamic environments. Consequently, we propose a more dynamic method that updates the Usage History to record the resource status within the cloud-native application during simulation runtime.
@@ 5.2 Usage History Updating
Based on the above observations, we attempt to quantify resource usage by the number and duration of cloudlets assigned to an instance for simulation. To address this, CloudNativeSim employs a linear approach to correlate the performance of instances with cloudlets’ resource usage. Using the specific calculation method, a cloudlet’s execution time depends on the CPU time allocated to it and the processing speed of the instance executing it. The execution time of cloudlets ultimately affects the calculation of response time. Similarly, the memory usage and bandwidth consumption of cloudlets are calculated using a similar approach. Once the resource consumption by the cloudlets is calculated, the corresponding instance’s usage data is updated and recorded in the usage history.
@@ 5.3 Service Scaling Policy
In the simulation of microservice-based architectures, the service scaling policy is a core function. It can leverage historical resource utilization data to maintain optimal performance, while preventing both resource wastage and inefficiency 27. During the operation of the simulator, a service scaling event is triggered at regular intervals to automatically check whether the current service requires scaling, such as when an instance’s CPU utilization consistently remains high 28. Once scaling is deemed necessary, the user-defined scaling policy is executed. In CloudNativeSim, we provide the most common horizontal and vertical scaling solutions, as outlined in Algorithms 5.3 and 5.3.
@@
Additionally, CloudNativeSim employs a threshold-based monitoring system that oversees both resource underutilization and overload scenarios. These thresholds are defined as ranges that trigger scaling decisions: scaling up occurs when resource utilization exceeds upper thresholds to maintain performance, while scaling down is initiated when utilization falls below lower thresholds to optimize resource efficiency.
Utilizing scaling policies greatly enhances the dynamism of CloudNativeSim, enabling it to flexibly respond to changes in workload, ensuring high service availability and efficient resource utilization. In Section 6.4, we will discuss the impacts of different scaling policies. Users can also customize auto-scaling policies, such as those based on response time or SLO violation rates 29, workload levels 30, and may even implement hybrid policies 31, 32.
In summary, CloudNativeSim implements three key scheduling policies: service placement, usage history tracking, and service scaling. These policies work together to provide a comprehensive solution for simulating cloud-native applications, with particular emphasis on resource efficiency and system adaptability. The extensible design allows users to customize these policies according to their specific requirements while maintaining the core functionality of the simulator.
@@ 6.1 Tests of Simulation Capacity
To test the simulation capacity of CloudNativeSim, we conducted a capacity test on a common PC with a 16-core, 2.2 GHz CPU, and limited the running memory to 4 GB. To evaluate the simulator’s extreme capacity, we considered the following parameters: the number of services, the total number of instances, the number of requests, and the number of cloudlets. VDifferent parameter configurations resulted varying running time on the realistic testbed (in seconds).
To ensure the reliability of our results, we designed 4 cases and tested 2 parameter sets for each case, conducting 5 repetitive tests for each set. Table 3 presents the average running time for these tests:
@@
To avoid randomness, we conducted ablation experiments in Case 1 and Case 2, selecting parameter sets with running time under 15 seconds for analysis. In Case 1, we limited the number of services to 1 and the number of instances to one thousand. Under these conditions, CloudNativeSim can simulate up to $10^{6}$ requests. In Case 2, we fixed the number of instances at 1 and the number of requests at $10^{4}$, demonstrating that CloudNativeSim can handle over $5\times 10^{4}$ services. Notably, the number of cloudlets also increases with the number of requests and services, leading to a substantial rise in the simulator’s internal running time.
@@ 6.2 Verification of Request Generation
To validate the performance of the Request Generator and the correctness of the modeling in Section 4.3.1, we select the most important parameters $(N_{c},v,p)$ to conduct four sets of test experiments on the Request Generator. In Section 4.3.1, Request Generation Modeling, the parameters $(N_{c},v,p)$ represent the total number of clients, the client growth rate, and the request inter-arrival time region, respectively. These parameters are crucial in determining the QPS and the evolution of the total number of clients within the simulation. The total number of clients $(N_{c})$ defines the upper limit of clients that can generate requests, directly impacting the load on the system. The client growth rate $(v)$ dictates how quickly new clients are added, influencing the rate at which the system load increases. The request inter-arrival time region $(p)$ determines the frequency of requests from each client, affecting the overall request rate. For example, QPS $(1000,100,[5,15])$ indicates a configuration with a maximum of 1000 clients, increasing 100 new clients per second, where each client generates requests with an inter-arrival time randomly distributed between 5 and 15 seconds. By collecting and plotting the growth curves of the number of clients, QPS, and total requests, we aimed to evaluate the performance of the Request Generator. Figure 10 illustrates the results produced by the Request Generator across four different test configurations.
@@
Based on the system capacity limits established in our previous experiments, we designed these test configurations to validate the Request Generator’s performance within the proven operational bounds. We selected three parameter configurations focusing on different client growth patterns. The first configuration (1000, 10) represents a gradual scaling scenario, where clients are added at a conservative rate of 10 per second until reaching the previously validated maximum of 1000 clients. The second configuration (1000, 100) maintains the same maximum client count but accelerates the growth rate to 100 clients per second, allowing us to examine how growth velocity affects request patterns while staying within known system limits. The third configuration (500, 100) keeps the rapid growth rate but operates at a more moderate client scale, providing insights into the Request Generator’s behavior under different steady-state loads. For request patterns, we tested two inter-arrival time regions: [3, 5] seconds for high-frequency requests and [5, 15] seconds for more moderate request rates. These combinations enable us to thoroughly evaluate the Request Generator’s performance across various realistic usage scenarios while operating within the validated system capacity established by our previous capacity testing experiments.
In Figure 10, the segmented curves in the three subplots exhibit distinct characteristics around 100 seconds, coinciding with the period of customer number growth, $\frac{N_{c}}{v}$. This period is highlighted due to its significance. During this time, the number of clients grows linearly at the spawn rate $v$, eventually stabilizing at the client number $N_{c}$. This process remains independent of the waiting interval $p$, which explains the presence of only three curves. This behavior aligns with the description provided in Equation 1. Initially, the QPS shows approximately linear growth before reaching an oscillatory convergence state. The speed of convergence is influenced by the spawn rate $v$, while the final converged value is determined by both $N_{c}$ and $p$. Specifically, the final value is negatively correlated with the value and interval size of $p$ and positively correlated with $N_{c}$, consistent with the description in Equation 3. In subplot 10(b), the blue and orange curves, which share the same $p$ and $N_{c}$, converge closely and oscillate around 100, as derived from Equation 3. For the total number of requests, the curve exhibits a nonlinear growth trend in the first segment and a linear growth trend in the second segment, consistent with the description in Equation 4.
@@ 6.3 Evaluation of Response Time Simulation
To assess the simulator’s capability in approximating the response time of actual applications, we conducted a stress test on the cluster using Locust, collecting average response time under varying numbers of users as the reference standard. In the Locust script, user wait time were set between 5 to 15 seconds, with a duration limit of 600 seconds. The same settings were applied in the simulator.
@@ 6.4 Demonstration of Scaling Policies
7 Conclusions, Discussions and Future Work
@@
7.1 Current Limitations
Although CloudNativeSim exhibits considerable potential in simulating cloud-native applications and environments, the current version has several inherent limitations:
  • Limited network traffic modeling. The current implementation lacks comprehensive network traffic simulation capabilities. While basic communication between microservices is supported, it does not fully capture complex network behaviors such as packet loss and bandwidth constraints in NetworkCloudSim 11, affecting the accuracy of performance predictions.
  • Lack of edge service mobility support. The current version does not support the mobility of edge services, crucial for simulating dynamic edge computing scenarios. Unlike iFogSim 12, 13, which effectively handles service mobility, current CloudNativeSim could not simulate scenarios where edge services need to migrate or adapt to moving edge devices.
  • Service discovery mechanism. The current implementation can be extended to include service discovery features developed in ServiceSim 15, which would enable more detailed modeling of service communication patterns in dynamic environments, particularly useful for large-scale deployments.
  • Limited fault tolerance simulation. The current implementation lacks comprehensive fault tolerance capabilities. While basic service failures can be simulated, it does not fully support advanced fault tolerance mechanisms, affecting the simulator’s ability to evaluate system reliability and resilience under various failure scenarios.
7.2 Future Work
Several promising directions can address the current limitations and enhance CloudNativeSim’s capabilities:
  • Implementation of service discovery and fault tolerance. Future work will focus on implementing comprehensive service discovery mechanisms and fault tolerance patterns, enabling more accurate simulation of service interactions in dynamic environments.
  • Enhanced monitoring and observability. Future versions will incorporate advanced monitoring capabilities, providing deeper insights into system behavior and performance bottlenecks, particularly crucial for large-scale deployments.
  • Optimizing large-scale scenario simulations. Future work will focus on implementing efficient memory management practices and improved service discovery mechanisms designed for large-scale simulations.
  • Advanced scheduling algorithms. Future development will incorporate more sophisticated scheduling algorithms to improve resource utilization and overall system performance.
  • Integration with container orchestration platforms. Future developments will include support for simulating container orchestration platforms like Kubernetes, enabling more realistic modeling of deployment strategies and resource management in cloud-native environments.
@@ 7.2 Future Work
Author Contributions
All authors participated in extensive discussions and contributed to the conceptualization, analysis, and finalization of this manuscript. The specific contributions are as follows: Minxian Xu conceived the original concept and provided overall guidance for the experimental methodology and manuscript development. Jingfeng Wu enhanced the initial concept, implemented the experimental framework, conducted comprehensive experiments, and authored the primary content of the manuscript. Yiyuan He focused on manuscript enhancement through detailed revision, language refinement, and improvement of the overall presentation quality. Kejiang Ye and Chengzhong Xu provided critical insights regarding application scenarios and scalability considerations, while also contributing to the manuscript’s refinement.
@@ Acknowledgments
This work is supported by Guangdong Basic and Applied Basic Research Foundation (No. 2024A1515010251, 2023B1515130002), Guangdong Special Support Plan (No. 2021TQ06X990), Shenzhen Basic Research Program under grants JCYJ20220818101610023, and JCYJ20240809180935001, Shenzhen Industrial Application Projects of undertaking the National key R & D Program of China (No. CJGJZD20210408091600002).